# The Democratization of Wealth Management: Hedged Mutual Fund Blockchain Protocol

### v1

@@
May 24, 2024
Keywords: Performance Fees; High Water Mark; Asset Price; Blockchain; Smart Contract; Investment Fund
@@ 2.1 Blockchain Investment Funds: Challenges and Solutions
There are several works that describe the application of blockchain technology to fund management and also to other aspects of financial intermediation (Manda et al., 2018; 2023; Patel, Migliavacca & Oriani 2022; Ren et al., 2023). These pioneering works provide a conceptual overview regarding the procedures pertaining to fund management that can be transferred to decentralized ledger based solutions. Streamlining some of the fund management processes onto blockchain can bring about significant operational efficiencies - that will reduce costs - and improved security due to the inherent cryptographic features of decentralized technologies - while making data publicly available for both reporting and decision making. (Fiergbor 2018; Ciriello 2021; Srivastava 2023) provide detailed examples pertaining to index funds - that is mutual funds and exchange traded funds (End-note 8).
@@
It is important to clarify that the existing works on blockchain fund management focus on providing managerial insights and high level overviews on how to approach the nuances of decentralized technology that can be applied to financial wealth management. In contract, we have designed a complete fund management blockchain protocol - including detailed mathematical formations and pointers - that is ready for software implementation.
@@
We have also created several innovations - related to investor protection and seamlessly levying fees - applicable to blockchain fund management in terms of adapting best practices from traditional finance to the blockchain realm. Initially, we describe the new techniques we have created at a high level so that it is easy to understand for a wide audience. Later sections give granularity steps - so that it helps participants understand the complexities involved and also to be able to build the necessary infrastructure. The outline of the various topics is given later in Section (2.3).
@@
Our novel techniques combine the superior mechanisms of investment funds - mutual funds and hedge funds - while provide detailed steps to implement them as blockchain smart contracts (Buterin 2014; Macrinici et al., 2017; Mohanta et al., 2018; Hu et al., 2019; Negara et al., 2021; Vacca et al., 2021; Sharma et al., 2023; End-note 9). We discuss how our designs overcome several blockchain bottlenecks and how we can make smart contracts smarter. Specifically, we illustrate how fund prices can be updated regularly like mutual funds and performance fees can be charged like hedge funds. In addition we show how mutually hedged blockchain investment funds can operate with investor protection measures such as high water marks and methodologies to offset trading related slippage costs when redemptions happen. We provide numerical illustrations of several scenarios related to the mechanisms we have tailored for blockchain implementation.
@@
The concepts we have developed for blockchain implementation can also be useful in traditional financial funds to calculate performance fees in a simplified manner. We show how blockchain technology can alleviate concerns regarding two issues we have highlighted with the operation of mutual funds and hedge funds. The ideas developed here illustrate on one hand, how blockchain can solve many issues faced by the traditional world and on the other hand, how many innovations from traditional finance can benefit decentralized finance and speed its adoption. This becomes an example of symbiosis between decentralized and traditional finance - bringing these two realms closer and breaking down barriers between such artificial distinctions - wherein the future will be about providing better risk adjusted wealth appreciation opportunities to end customers through secure, reliable, accessible and transparent services - without getting too caught up about how such services are being rendered.
@@ 2.2 Fees for Protection with Appreciation: Quid Pro Quo
The community driven spirit of DeFi makes it essential to have some form of profit sharing, wherein part of any excess fees generated is given back to long term investors of the project (Singh & Kim 2019; Wang et al., 2019; El Faqir et al., 2020; Wu et al., 2021; Zheng & Boh 2021; Ballandies et al., 2022; Liu et al., 2022; Kitzler et al., 2023; Saurabh et al., 2023). Fueling the decentralized ethos of profit sharing are many grievances related to high fees in the traditional financial sector with counter arguments being provided about competition between funds restricting fees to reasonable levels (Golec 1996; Coates & Hubbard 2007; Khorana et al., 2009; Malkiel 2013; Philippon 2015; Feldman et al., 2020). With blockchain technology it becomes easy to assess profits, fees and related metrics - due to the transparency of operations across all types of decentralized investment vehicles - given that making information publicly available is a foundational principle built right into the genes.
@@
While the importance of sharing the proceeds - profits from operations – with all the participants cannot be emphasized enough, it is equally - if not more - important to ensure that the community understand the need to have funds set aside for a rainy day and the necessity of being able to invest in projects that can prepare the community for the a better tomorrow. Most protocols that wish to follow the mechanisms designed here need to emphasize to investors - community - that the performance fees are to ensure that there are enough funds to sustain operating expenses and fund future growth plans. The performance fees will ensure that investments remains stable across market cycles - and continue to grow - rewarding investors over the long term. Additional performance fees above a threshold will be the returned back to the community (Kashyap 2021).
@@
An extremely popular investor protection mechanism in the traditional finance world is the idea of a high water mark (HWM: Goetzmann, et al., 2003; Guasoni & Obłój 2016; End-note 10). The simple summary of this concept is that performance fees are charged only when investors are entitled to a profit derived from their original principal. This is perhaps best clarified with a simple numerical illustration.
@@
For example, let us say an investor deposits 10,000 USD. After some time, the invested amount grows to 14,000 USD, at which a high water mark is established. The profit in this case is 4,000 USD. A part of this profit is taken as performance fees. After this, if the value of the investment goes down to say 12,000 no performance fees are charged until the value of investment climbs back above 14,000, the high water mark. The bottomline is that unless a tangible wealth increase is generated for every investor - at a holistic level - no performance fees are paid. This creates a strong incentive for the team to produce solid - and quantifiable measurable - returns for the investors.
@@
This simple scenario can get extremely complicated when there are multiple investors who deposit at different levels of the fund price. Tracking all this in a smart contract - with the current state of blockchain technology - is extremely hard and can be deemed almost impossible (Giancaspro 2017; Wang, et al., 2018; Huang et a., 2019; Zou, et al., 2019; Sayeed et al., 2020; Zheng, et al., 2020; Tolmach et al., 2021; Kannengießer et al., 2021). To be able to accommodate these complexities we have found a novel solution that works elegantly - is rather straightforward to implement as a smart contract - and provides the same level of protection to every single investor. Our solution - which utilizes weighed average calculations - is also mathematically identical - in terms of aggregate fees and proceeds - to what investment funds in the traditional world have been doing for decades.
@@
In addition, the detailed algorithm we have developed - to create a fund management protocol - ensures that other investor protection schemes are incorporated. In particular we have created techniques to maintain fairness while investors enter or exit the fund. Limits - maximum amounts - are set on the total amounts that can enter or leave the fund during any time interval. The overall objective is to reward investor loyalty and their preference for the staying longer with the fund, which will benefit all participants - the entire community.
@@
Also, the buying and selling of assets affects asset prices and is known as slippage (Bertsimas & Lo 1998; Kashyap 2020; End-note 12). We distribute the price appreciation - and the depreciation - when investor money enters - or leaves - the fund across all participants using the treasury - to finance shortfalls or to collect surplus amounts - so that no one benefits from large deposits or gets penalized when large sells are made.
@@ 2.3 Outline of the Sections Arranged Inline
Section (3) gives an algorithm to accept deposits and withdrawal requests from investors after calculating the price of fund shares or tokens. Section (4) is a discussion of a novel technique to charge performance fees, and maintain high water marks, despite the limitations of smart contracts. Section (4.1) considers the challenges of performance fee calculations in a decentralized environment. Section (4.2) outline issues that funds face when the NAV - due to falling market prices - trends lower than the fund water mark for charging fees. Section (4.3) outlines our solutions to address the problem in Section (4.2) tailored to perform seamlessly on blockchain networks. Section (4.4) is a discussion of how fee levies can happen in the absence of our solutions in Section (4.3) and to help readers understand the significance of the techniques we have developed.
@@ 3 Periodic Blockchain Fund Management Methods: Sequences of Steps
The following algorithm captures the sequence of steps that need to be carried out at periodic intervals to accomplish secure fund management entirely on a blockchain environment. The interaction of the various processes described in the following points facilitate secure movement of assets, rigorous risk management and rebalancing the portfolio to adhere to asset weights that match risk and return objectives.
@@
    Check Existing Number of Tokens for the Alpha investment fund (across all chains). Tokens represent the number of units or the quantity of any asset. They are the same as shares of stocks in the traditional financial world.
@@
    Aggregate the deposit amounts and withdrawal amounts so that the net total amount received for investment or for redemption is less than the corresponding maximum amount.
@@
    • The maximum amounts for deposit or withdraw for any rebalancing event is to ensure that there is a limit to how much funds can be taken out or invested into the fund during any given time interval.
@@
    • We have to look at two cases depending on whether money will flow into the fund or out of the fund during this rebalancing sequence of events. That is we set flags $NETDPSTIND_{t}$ or $NETWDRWIND_{t}$ which will indicate whether the net amount is positive (net deposit) or negative (net withdrawal).
@@
    • $W_{t}$ is the total number of withdrawals requested at time $t$ before this particular rebalance sequence of events has started.
    • $D_{t}$ is the total number of deposit requests at time $t$ before this particular rebalance sequence of events has started.
@@
        If $RBLNCSLIPPAGE_{t}>0$ is positive, send this amount to the treasury and burn or remove Alpha tokens equal to this amount from the treasury using the NAV or Alpha Price calculated in Step (6).
@@
        If $RBLNCSLIPPAGE_{t}<0$ is negative, take stable coins (Ante, Fiedler & Strehle 2021; End-note 13) from the treasury equal to this amount and issue new Alpha tokens equal to this amount using the NAV or Alpha Price calculated in Step (6). Send the tokens to the treasury.
@@
        • This is because the total number of tokens now includes the tokens burned or minted for the amount, $RBLNCSLIPPAGE_{t}$, that were removed or issued in Sub-Steps (7(b)iv; 7(b)v).
@@
        Remove (burn) Alpha tokens to fulfill the withdraw requests that were accepted for redemption and issue (mint) new Alpha Tokens to fulfill the deposit requests that were accepted for investment as necessary using the NAV or Alpha Price calculated in Step (6).
@@ 4 Automated Performance Fee Levy Techniques
In any wealth management fund, to calculate performance fees properly, all the transactions done by all investors have to be recorded and utilized when we perform the corresponding computations to levy the fees. To give an example, an investor could be holding 10000 tokens of the Alpha fund in his wallet, but he could have done several buy and sell transactions - at different prices - to arrive at this overall position. That is, a position for an investor is comprised of multiple transactions. Taking into account the transactions from different investors implies a large amount of transaction level data.
@@
In traditional hedge funds, accountants look at all investor positions - also the transactions as necessary - every quarter - or other such infrequent intervals - and calculate how much performance fees are due. Most hedge funds allow new investments - or redemptions - only at certain times. This fund entry and exit is usually not very often, perhaps at quarterly or larger durations. Hence the tasks of the fund accountants are relatively straightforward and they do performance fee calculations easily using specialized software or even using excel based tools.
@@ 4.1 Blockchain Bottlenecks and Making Smart Contracts Smarter
There are two broad approaches to blockchain performance fee calculations. They can be done off-chain or on-chain. Off-chain computations refer to using software routines or calculations outside the blockchain. On-chain calculations are done entirely within the blockchain environment. As we know from the foundations of computing science, if we store a lot of data in the computer memory we can reduce the data we have to store external to the computer on databases or other file based storage systems. If we do not store data within the computer, we have to expend resources reading and writing data to external infrastructure. Also, the more data we have, the more intensive the calculations or the computation time of the corresponding algorithms increases with increasing data size. Hence, we need more sophisticated or complex algorithms when we have to handle large data-sets (Knuth 1973; 1997; Aho & Hopcroft 1974; Horowitz & Sahni 1982; Aho 2012).

### v2

@@
July 29, 2024
Keywords: Performance Fees; High Water Mark; Asset Price; Blockchain; Smart Contract; Investor Protection; Investment Fund
@@ 2.1 Blockchain Investment Funds: Challenges and Solutions
There are several works that describe the application of blockchain technology to fund management and also to other aspects of financial intermediation (Manda et al., 2018; 2023; Patel, Migliavacca & Oriani 2022; Ren et al., 2023). These pioneering works provide a conceptual overview regarding the procedures pertaining to fund management that can be transferred to decentralized ledger based solutions. Streamlining some of the fund management processes onto blockchain can bring about significant operational efficiencies - that will reduce costs - and improved security due to the inherent cryptographic features of decentralized technologies, while making data publicly available for both reporting and decision making. (Fiergbor 2018; Ciriello 2021; Srivastava 2023) provide detailed examples pertaining to index funds - that is mutual funds and exchange traded funds (End-note 8).
@@
It is important to clarify that the existing works on blockchain fund management focus on providing managerial insights and high level overviews on how to approach the nuances of decentralized technology that can be applied to financial wealth management. In contrast, we have designed a complete fund management blockchain protocol - including detailed mathematical formulations and pointers - that is ready for software implementation.
@@
We have also created several innovations - related to investor protection and seamlessly levying fees - applicable to blockchain fund management in terms of adapting best practices from traditional finance to the blockchain realm. Initially, we describe the new techniques we have created at a high level, so that they are easy to understand for a wide audience. Later sections give granular steps - so that they can help participants comprehend the complexities involved and also to be able to build the necessary infrastructure. The outline of the various topics is given later in Section (2.3).
@@
Our novel techniques combine the superior mechanisms of investment funds - mutual funds and hedge funds - while provide detailed steps to implement them as blockchain smart contracts (Buterin 2014; Macrinici et al., 2017; Mohanta et al., 2018; Hu et al., 2021; Negara et al., 2021; Vacca et al., 2021; Sharma et al., 2023; End-note 9). We discuss how our designs overcome several blockchain bottlenecks and how we can make smart contracts smarter. Specifically, we illustrate how fund prices can be updated regularly like mutual funds and performance fees can be charged like hedge funds. In addition, we show how mutually hedged blockchain investment funds can operate with investor protection measures such as high water marks and methodologies to offset trading related slippage costs when redemptions happen. We provide numerical illustrations of several scenarios related to the mechanisms we have tailored for blockchain implementation.
@@
The concepts we have developed for blockchain implementation can also be useful in traditional financial funds to calculate performance fees in a simplified manner. We show how blockchain technology can alleviate concerns regarding two issues we have highlighted with the operation of mutual funds and hedge funds. The ideas developed here illustrate, on one hand, how blockchain can solve many issues faced by the traditional world, and on the other hand, how many innovations from traditional finance - tried and tested over time - can benefit decentralized finance and speed its adoption. This becomes an example of symbiosis between decentralized and traditional finance - bringing these two realms closer and breaking down barriers between such artificial distinctions - wherein the future will be about providing better risk adjusted wealth appreciation opportunities to end-customers through secure, reliable, accessible and transparent services - without getting too caught up about how such services are being rendered.
@@ 2.2 Fees for Protection with Appreciation: Quid Pro Quo
The community driven spirit of DeFi makes it essential to have some form of profit sharing, wherein part of any excess fees generated is given back to long term investors of the project (Singh & Kim 2019; Wang et al., 2019; El Faqir et al., 2020; Wu et al., 2021; Zheng & Boh 2021; Ballandies et al., 2022; Liu et al., 2022; Kitzler et al., 2023; Saurabh et al., 2023). Fueling the decentralized ethos of profit sharing are many grievances related to high fees in the traditional financial sector, with counter arguments being provided about competition between funds restricting fees to reasonable levels (Golec 1996; Coates & Hubbard 2007; Khorana et al., 2009; Malkiel 2013; Philippon 2015; Feldman et al., 2020). With blockchain technology it becomes easy to assess profits, fees and related metrics - due to the transparency of operations across all types of decentralized investment vehicles - given that making information publicly available is a foundational principle built right into the genes.
@@
While the importance of sharing the proceeds - profits from operations – with all the participants cannot be emphasized enough, it is equally - if not more - important to ensure that the community understands the need to have funds set aside for a rainy day and the necessity of being able to invest in projects that can prepare the community for a better tomorrow. Most protocols that wish to follow the mechanisms designed here need to emphasize to investors - community - that the performance fees are to ensure that there are enough funds to sustain operating expenses and fund future growth plans. The performance fees will ensure that investments remains stable across market cycles - and continue to grow - rewarding investors over the long term. Additional performance fees, above a threshold, will be returned back to the community (Kashyap 2021).
@@
An extremely popular investor protection mechanism, in the traditional finance world, is the idea of a high water mark (HWM: Goetzmann, et al., 2003; Guasoni & Obłój 2016; End-note 10). The simple summary of this concept is that performance fees are charged only when investors are entitled to a profit derived from their original principal. This is, perhaps, best clarified with a simple numerical illustration.
@@
For example, let us say an investor deposits 10,000 USD. After some time, the invested amount grows to 14,000 USD, at which a high water mark is established. The profit in this case is 4,000 USD. A part of this profit is taken as performance fees. After this, if the value of the investment goes down to say 12,000 no performance fees are charged until the value of investment climbs back above 14,000, the high water mark. The bottomline is that unless a tangible wealth increase is generated for every investor - at a holistic level - no performance fees are paid. This creates a strong incentive for the team to produce solid - and quantifiably measurable - returns for the investors.
@@
This simple scenario can get extremely complicated when there are multiple investors who deposit at different levels of the fund price. Tracking all this in a smart contract - with the current state of blockchain technology - is extremely hard and can be deemed almost impossible (Giancaspro 2017; Wang, et al., 2018; Huang et a., 2019; Zou, et al., 2019; Sayeed et al., 2020; Zheng, et al., 2020; Tolmach et al., 2021; Kannengießer et al., 2021). To be able to accommodate these complexities we have found a novel solution that works elegantly - is rather straightforward to implement as a smart contract - and provides the same level of protection to every single investor. Our solution - which utilizes weighed average calculations - is also mathematically identical, in terms of aggregate fees and proceeds, to what investment funds in the traditional world have been doing for decades.
@@
In addition, the detailed algorithm we have developed - to create a fund management protocol - ensures that other investor protection schemes are incorporated. In particular, we have created techniques to maintain fairness while investors enter or exit the fund. Limits - maximum amounts - are set on the total amounts that can enter or leave the fund during any time interval. The overall objective is to reward investor loyalty and their preference for staying longer with the fund, which will benefit all participants - the entire community.
@@
Also, the buying and selling of assets affects asset prices and is known as slippage (Bertsimas & Lo 1998; Kashyap 2020; End-note 12). We distribute the price appreciation, and the depreciation, when investor money enters, or leaves, the fund across all participants using the treasury - to finance shortfalls, or to collect surplus amounts - so that no one benefits from large deposits, or gets penalized when large sells are made.
@@ 2.3 Outline of the Sections Arranged Inline
Section (3) gives an algorithm to accept deposits and withdrawal requests from investors, after calculating the price of fund shares or tokens. Section (4) is a discussion of a novel technique to charge performance fees, and maintain high water marks, despite the limitations of smart contracts. Section (4.1) considers the challenges of performance fee calculations in a decentralized environment. Section (4.2) outlines issues, for charging fees, that funds face when the NAV trends lower - due to falling market prices - than the fund water mark. Section (4.3) outlines our solutions to address the problem in Section (4.2) tailored to perform seamlessly on blockchain networks. Section (4.4) is a discussion of how fee levies can happen in the absence of our solutions in Section (4.3) and to help readers understand the significance of the techniques we have developed.
@@ 3 Periodic Blockchain Fund Management Methods: Sequences of Steps
The following algorithm captures the sequence of steps, that need to be carried out at periodic intervals, to accomplish secure fund management entirely on a blockchain environment. The interaction of the various processes described in the following points facilitate secure movement of assets, rigorous risk management and rebalancing the portfolio to adhere to asset weights that match risk and return objectives.
@@
    Check Existing Number of Tokens for the Alpha investment fund (across all chains). Tokens represent the number of units, or the quantity of any asset. They are the same as shares of stocks in the traditional financial world.
@@
    Aggregate the deposit amounts and withdrawal amounts so that the net total amount received for investment, or for redemption, is less than the corresponding maximum amount.
@@
    • The maximum amounts for deposit, or withdraw, for any rebalancing event is to ensure that there is a limit to how much funds can be taken out of, or invested into, the fund during any given time interval.
@@
    • We have to look at two cases depending on whether money will flow into the fund, or out of the fund, during this rebalancing sequence of events. That is we set flags $NETDPSTIND_{t}$ or $NETWDRWIND_{t}$, which will indicate whether the net amount is positive (net deposit) or negative (net withdrawal).
@@
    • $W_{t}$ is the total number of withdrawals requested at time $t$, before this particular rebalance sequence of events has started. It is to be understood henceforth, unless otherwise stated, that the aggregations include items since the last rebalance event has completed.
    • $D_{t}$ is the total number of deposit requests at time $t$, before this particular rebalance sequence of events has started.
@@
        If $RBLNCSLIPPAGE_{t}>0$ is positive, send this amount to the treasury and burn, or remove, Alpha tokens equal to this amount from the treasury using the NAV, or Alpha Price, calculated in Step (6).
@@
        If $RBLNCSLIPPAGE_{t}<0$ is negative, take stable coins (Ante, Fiedler & Strehle 2021; End-note 13) from the treasury equal to this amount and issue new Alpha tokens equal to this amount using the NAV, or Alpha Price, calculated in Step (6). Send the tokens to the treasury.
@@
        • This is because the total number of tokens now includes the tokens burned, or minted, for the amount, $RBLNCSLIPPAGE_{t}$, that were removed, or issued, in Sub-Steps (7(b)iv; 7(b)v).
@@
        Remove (burn) Alpha tokens to fulfill the withdraw requests that were accepted for redemption and issue (mint) new Alpha Tokens to fulfill the deposit requests that were accepted for investment as necessary using the NAV, or Alpha Price, calculated in Step (6).
@@ 4 Automated Performance Fee Levy Techniques
In any wealth management fund, to calculate performance fees properly, all the transactions done by all investors have to be recorded and utilized when we perform the corresponding computations to levy the fees. To give an example, an investor could be holding 10,000 tokens of the Alpha fund in his wallet, but he could have done several buy and sell transactions - at different prices - to arrive at this overall position. That is, a position for an investor is comprised of multiple transactions. Taking into account the transactions from different investors implies a large amount of transaction level data.
@@
In traditional hedge funds, accountants look at all investor positions - also the transactions as necessary - every quarter - or other such infrequent intervals - and calculate how much performance fees are due. Most hedge funds allow new investments - or redemptions - only at certain times. This fund entry and exit is usually not very often, perhaps at quarterly, or longer durations. Hence the tasks of the fund accountants are relatively straightforward and they do performance fee calculations easily using specialized software, or even using excel based tools.
@@ 4.1 Blockchain Bottlenecks and Making Smart Contracts Smarter
There are two broad approaches to blockchain performance fee calculations. They can be done off-chain or on-chain. Off-chain computations refer to using software routines or calculations outside the blockchain. On-chain calculations are done entirely within the blockchain environment. As we know from the foundations of computing science, if we store a lot of data in the computer memory we can reduce the data we have to store external to the computer on databases, or other file based storage systems. If we do not store data within the computer, we have to expend resources reading and writing data to external infrastructure. Also, the more data we have, the more intensive the calculations, or the computation time of the corresponding algorithms increases with increasing data size. Hence, we need more sophisticated, or complex, algorithms when we have to handle large data-sets (Knuth 1973; 1997; Aho & Hopcroft 1974; Horowitz & Sahni 1982; Aho 2012).